OpenStudy (anonymous):

Given the equation 6xsquared-y=98 , find x if y = –2. A.16 B.10 C.4 D.60

OpenStudy (fifciol):

\[6x^2-(-2)=98\] \[6x^2+2=98\] \[6x^2=96 /:6\] \[x^2=16 \rightarrow x=4 \] or x= -4 ans C

OpenStudy (anonymous):

thanks :)